OpenStudy (anonymous):

What is the slope of the line containing the points (2,5) and (-3,5)? 1. -5 2. 5 3. undefined 4. 0

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ 5}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ 5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (anonymous):

Did you use the formula (x1,y1) (x2, y2)

OpenStudy (anonymous):

Use that formula^

OpenStudy (anonymous):

Y = -2x -1

OpenStudy (anonymous):

(x1, y1) = (2, 5) and (x2, y2)= (-3, 5)

OpenStudy (anonymous):

I got -1.3?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Don't know

OpenStudy (anonymous):

Hmm

OpenStudy (jdoe0001):

so... what did you get for the slope?

OpenStudy (anonymous):

If you don't know how to solve slopes then you should go here for future help. http://www.ehow.com/how_2362821_find-slopes.html

OpenStudy (igreen):

\(m = \dfrac{y_2-y_1}{x_2-x_1}\) This is the slope formula. In this case: \(y_2 = 5\) \(y_1 = 5\) \(x_2 = -3\) \(x_1 = 2\) Plug them in: \(m = \dfrac{5-5}{-3-2}\) 5 - 5 cancels out leaving you with -3 - 2, and what's that? @Daniella33

OpenStudy (igreen):

What's -3 - 2? @Daniella33

OpenStudy (anonymous):

Do you have a graphing calculator, Daniella?